The amounts of soft drink machine is designed to dispense for each drink are normally distributed, with a mean
of 11.8 fluid ounces and a standard deviation of 0.3 fluid ounce. A drink is randomly selected.
Find the probability that the drink is more than 12.2 fluid ounces. Can this be considered an unusual event?
Explain your reasoning.
Answer #1

Solution :

Given that ,

mean = = 11.8

standard deviation = = 0.3

P(x > 12.2) = 1 - p( x< 12.2)

=1- p P[(x - ) / < (12.2 - 11.8) / 0.3]

=1- P(z < 1.33)

Using z table,

= 1 - 0.9082

= 0.0918

No, its not considered unusual, because probability more than 5%